OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (nurali):

C = coach; T = Tim C = T + 10; C-3 = 2(T-3), so C = 2T - 6 + 3 = 2t - 3 2t - 3 = t + 10 t = 13, Tim's age now, so coach is 23; three years ago they were 10 & 20

OpenStudy (nurali):

if correct please click best response.

OpenStudy (anonymous):

could you explain because im not getting it

OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Let's call C the coach's current age, and T tim's current age, right now the coach is ten years older than tim, which means that: C = T + 10 Three years ago, coach johnson was twice tims age (C - 3 is the age of the couch 3 years earlier, and T - 3 tim's age 3 years earlier), C - 3 = (T - 3) * 2 Now we have 2 equations with two unknows that you can solve: C = T + 10 C = (T - 3)*2 + 3 Now this means that (T - 3)*2 + 3 = T + 10 2T - 6 + 3 = T + 10 2T = T + 13 T = 13 So we found tim's current age, which is 13 years old